OpenStudy (anonymous):

4 times the lesser of the two consecutive even intergers is 12 less than twice the greater #'s . what are the intergers, and how did you get them?

hero (hero):

consecutive even numbers are always expressed as x + (x + 2) where x = the first number and x + 2 = the second number. 4x = 2(x+2) -12 4x = 2x + 4 - 12 4x - 2x = -12 2x = -8 x = -4 x + 2 = -4 + 2 = -2 To check replace x above: 4x = 2(x+2) - 12 4(-4) = 2(-4+2) -12 -16 = 2(-2) - 12 -16 = -4 - 12 -16 = -16

hero (hero):

Somehow, you still got -4

OpenStudy (radar):

But I failed to read the question and solve for even integers.

hero (hero):

It happens. I've learned to read the question twice and take my time.

OpenStudy (radar):

I had -5 and -4, just lucked out o the -4 lol

OpenStudy (anonymous):

thanks.